# MetaFold: Language-Guided Multi-Category Garment Folding Framework via Trajectory Generation and Foundation Model

### v1

@@ Abstract
Garment folding is a common yet challenging task in robotic manipulation. The deformability of garments leads to a vast state space and complex dynamics, which complicates precise and fine-grained manipulation. Previous approaches often rely on predefined key points or demonstrations, limiting their generalization across diverse garment categories. This paper presents a framework, MetaFold, that disentangles task planning from action prediction, learning each independently to enhance model generalization. It employs language-guided point cloud trajectory generation for task planning and a low-level foundation model for action prediction. This structure facilitates multi-category learning, enabling the model to adapt flexibly to various user instructions and folding tasks. Experimental results demonstrate the superiority of our proposed framework. Supplementary materials are available on our website: https://meta-fold.github.io/.
@@ I Introduction
One of the major challenges in general robotic applications is the manipulation of deformable objects [1]. Tasks that humans perform intuitively, such as folding clothes, present significant challenges for robots. This difficulty arises because deformable objects exhibit diverse forms of deformation, resulting in large state space and complex dynamics [2]. These factors complicate the prediction of future configurations, the planning of manipulation points, and the precise control of force and positioning.
@@
Predicting future states and actions from arbitrary initial states introduces highly complex dynamics, posing significant challenges to the learning process. By contrast, decomposing the task and focusing specifically on predicting garment states during the folding process simplifies the learning task and further facilitates multi-category generalization, benefiting from the strong capabilities of generative models.
@@
Accurate prediction of future states in garment manipulation requires an effective representation of garment configurations. Point cloud trajectories fulfill this need by providing a comprehensive spatial model that captures any point in space and encodes temporal changes in object states. However, existing methods often rely on structural representations, such as key points or skeletal models [3], and depend on demonstrations or heuristic-based policies [4, 5, 6, 7, 8]. This dependence restricts these models to specific object configurations, limiting their adaptability across garment categories. In contrast, by using point cloud trajectories instead of structural representations, we maintain spatial and temporal awareness of the states, enabling more precise garment folding. Additionally, employing generative modeling to create point cloud trajectories facilitates multi-category learning and generalization within a unified framework, yielding robust performance across diverse garment manipulation tasks.
@@
  • We introduce a novel framework for garment manipulation that disentangles state planning from action prediction. This hierarchical separation enables the model to focus on distinct components, thereby reducing training complexity.
@@
  • We developed the first point cloud trajectory dataset for folding garments in multiple categories, accompanied by corresponding language descriptions as guidance.
@@ II-A Garment Folding
Garment folding is a challenging task due to the deformable and high-dimensional nature of fabrics, which has been extensively studied in the literature [12, 13, 14, 15, 16].
@@
Cloth Funnels [8] trained robots to align garments into predefined deformable configurations, simplifying the manipulation process for downstream folding. UniGarmentManip [5] explored learning dense visual correspondences for garment manipulation with category-level generalization. FabricFlowNet [20] developed a cloth manipulation policy that leverages flow as both input and action representation to enhance performance.  [21] and [22] have proposed learning cloth dynamics to improve manipulation capabilities.
@@ II-B Language-Conditioned Object Manipulation
Language-conditioned manipulation has garnered significant attention due to its user-friendly interaction and strong generalization capabilities [24, 25, 26, 27, 28, 29]. [30] introduced a Transformer-based model for mapping natural language commands to robot actions in 3D tasks. CLIPort[31] combines CLIP’s semantic understanding with Transporter’s spatial precision for language-guided manipulation.
With recent advances in Large Language Models (LLMs), researchers have explored leveraging LLMs to integrate language understanding into cloth folding. GPT-Fabric [10] employs GPT to directly output actions, informing a robot where to grasp and pull a simple square-shaped fabric.  [32] introduced a method for language-conditioned manipulation that utilizes graph dynamics to improve the understanding and prediction of deformable object behaviors.
@@ III Problem Formulation
The goal of language-guided garment folding is to generate a sequence of actions $\{\boldsymbol{a}_{i}\}_{i=1}^{n}$ to fold the garment to a goal point cloud configuration $\mathcal{P}_{goal}$, given the 3D point cloud observation $\mathcal{P}\in\mathbb{R}^{N\times 3}$ with $N$ points and the language guidance $\mathcal{L}$.
@@
(1) Given the current $\mathcal{P}$ with $N$ points and $\mathcal{L}$, the goal of the point cloud trajectory generation task is to generate the trajectory $\mathcal{T}=\{\mathcal{P}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times N\times 3}$ that represents the evolution of the point cloud over time, where $M$ is the number of frames.
(2) Given two point clouds $(\mathcal{P},\mathcal{P^{\prime}})$, the aim of the ManiFoundation [11] model is to predict the action $\boldsymbol{a}=\{c_{i}\}_{i=1}^{n}$, which is defined as a set of contact syntheses. The contact synthesis for end effector $i$ is $c_{i}=(\boldsymbol{p},\boldsymbol{s})$, where $\boldsymbol{p}\in\mathbb{R}^{3}$ is the contact position and $\boldsymbol{s}\in\mathbb{R}^{3}$ is the corresponding grasping distance based on the trajectory. We slice the generated garment point cloud trajectory $\mathcal{T}$ and input the segments into the ManiFoundation model to predict an action $\boldsymbol{a}=\mathcal{M}_{MF}(\mathcal{P},\mathcal{P^{\prime}})$, where $\{\mathcal{P},\mathcal{P^{\prime}}\}\subseteq\mathcal{T}$.
(3) Afterward, the robot executes the action $\boldsymbol{a}$, manipulating the garment to a new configuration $\mathcal{P}^{\prime}=\mathcal{M}_{Robot}(\mathcal{P},\boldsymbol{a})$.
@@ IV-A Overview
Our framework is illustrated in Figure 2. In this section, we provide a detailed explanation of language-guided point cloud trajectory generation (Sec. IV-B) and closed-loop manipulation (Sec. IV-C). Below, we first introduce training and inference pipelines.
@@ IV-B2 Conditional Point Cloud Trajectory Generation
As discussed earlier, point cloud trajectories can model state transitions, allowing us to control intermediate states during garment folding. To achieve this, we design a point cloud trajectory generation model. This approach enables a comprehensive generation and control of garment movement over time, providing a complete understanding of its dynamics. From a spatial perspective, this method tracks all points in the initial frame and maps the subsequent trajectory of each point. Temporally, each frame represents the garment’s state at a specific moment.
@@
In training, both the spatial features of the point cloud $\mathcal{F}_{\mathcal{P}}$ and the language features $\mathcal{F}_{\mathcal{L}}$ are combined with the ground truth trajectory and used as conditional inputs to the CVAE encoder. This optimizes the distribution of the latent variable $z$. In both training and inference, samples are drawn from the distribution of $z$ and combined with the spatial features of the point cloud $\mathcal{F}_{\mathcal{P}}$. The language features $\mathcal{F}_{\mathcal{L}}$ as conditional inputs to the CVAE decoder to generate trajectory features $\mathcal{F}_{\mathcal{T}}\in\mathbb{R}^{M\times N\times 256}$. An additional MLP is used to reduce the dimensionality of the output, providing the generated trajectory $\mathcal{T}\in\mathbb{R}^{M\times N\times 3}$ for the current frame. The Visualization of generated trajectories is in Figure 4.
@@ IV-B3 Semantic Generalization
We utilize a pre-trained large language model (Meta-Llama-3.1-8B-Instruct) [45] to extract semantic features. Specifically, we predefine a set of garment folding descriptions to strengthen the encoding process. During inference, we also employ LLaMA as the foundational language model to process user-provided guidance. The model matches user input to our predefined descriptions, effectively reducing the complexity of training on diverse language instructions by constraining conditions within a defined range. This approach leverages the large language model to achieve open-vocabulary generalization.
@@ IV-C2 ManiFoundation Model
Additionally, the predictions from the ManiFoundation model are influenced by random seeds. To mitigate this variability, we employ a model ensemble approach using 160 different random seeds to generate multiple predictions. When the distance between two prediction results is less than $\varepsilon$, they are considered similar and grouped. If a prediction result does not match any existing group, a new group is created. We compute the arithmetic mean of the point positions within a group and select the point closest to this average, along with its corresponding force, as the output contact point and force direction.
@@ V-C Baselines
For the garment folding task (Q1), we use UniGarmentManip [5] as the baseline. This method, previously demonstrated state-of-the-art (SOTA) performance in individual garment folding tasks [5], supports folding various garment types, thereby enabling a comprehensive comparison of folding performance across different categories. We used GPT-Fabric [10] as a baseline, which is a recent method that leverages large language models to select key points.
@@ V-D Results and Analysis
Although our approach is zero-shot on the CLOTH3D dataset, we achieve performance that matches or even surpasses the baseline, demonstrating the strong generalization of our method to a wide range of new garments. In garment folding, our model tends to fold clothing into a smaller and more compact form, which aligns well with human expectations for neatness and compactness. In general, our approach outperforms the baselines in most tasks and metrics. Therefore, we can address Q1, demonstrating our superior performance in garment folding.
@@
Furthermore, our model supports folding based on user-provided language inputs, even if the folding sequence (Down $\rightarrow$ Left $\rightarrow$ Right) has not appeared in the training data. Figure 5 illustrates the folding results of our model under different language instructions.
@@ V-F Real-World Experiments
To obtain the garment point cloud, we first segmented the RGB image captured by an RGB-D camera using SAM2 [51] to generate a garment mask. This mask was then used to filter the corresponding depth data, extracting the actual point cloud of the clothing. The resulting point cloud served as the input for our algorithm. The results of real-world experiments are demonstrated in Fig. 6.
@@ VI Conclusion
In this work, we proposed a comprehensive framework, MetaFold, for garment folding that supports multi-category folding tasks guided by user language instructions. This framework adopts a disentangled structure consisting of a point cloud trajectory generation model and a low-level action prediction model, utilizing closed-loop control for effective garment manipulation. We also constructed a point cloud trajectory dataset for garment folding, encompassing various folding methods across different garment types. Experimental results demonstrate that our approach achieves state-of-the-art performance in multi-category and language-guided garment folding tasks. We believe that MetaFold represents a significant step forward in applying trajectory generation to deformable object manipulation, marking an important milestone toward advancing spatial intelligence.

### v2

@@ Abstract
Garment folding is a common yet challenging task in robotic manipulation. The deformability of garments leads to a vast state space and complex dynamics, which complicates precise and fine-grained manipulation. In this paper, we present MetaFold, a unified framework that disentangles task planning from action prediction and learns each independently to enhance model generalization. It employs language-guided point cloud trajectory generation for task planning and a low-level foundation model for action prediction. This structure facilitates multi-category learning, enabling the model to adapt flexibly to various user instructions and folding tasks. We also construct a large‑scale MetaFold dataset comprising folding point cloud trajectories for a total of 1210 garments across multiple categories, each paired with corresponding language annotations. Extensive experiments demonstrate the superiority of our proposed framework. Supplementary materials are available on our website: https://meta-fold.github.io/.
@@ I Introduction
One of the major challenges in general robotic applications is the manipulation of deformable objects [1]. Tasks that humans perform intuitively, such as folding clothes, present significant challenges for robots. This difficulty arises because deformable objects exhibit diverse forms of deformation, resulting in a large state space and complex dynamics [2]. These factors complicate the prediction of future configurations, the planning of manipulation points, and the precise control of force and positioning.
@@
Predicting future states and actions from arbitrary initial states introduces highly complex dynamics, posing significant challenges to the learning process. By contrast, decomposing the task and focusing specifically on predicting garment states during the folding process simplifies the learning process and further facilitates multi-category generalization, benefiting from the strong capabilities of generative models.
@@
Accurate prediction of future states in garment manipulation requires an effective representation of garment configurations. Point cloud trajectories fulfill this need by providing a comprehensive spatial model that captures any point in space and encodes temporal changes in object states. However, existing methods often rely on structural representations, such as key points or skeletal models [3], and depend on demonstrations or heuristic-based policies [4, 5, 6, 7, 8]. This dependence restricts these models to specific object configurations, limiting their adaptability across garment categories. In contrast, by using point cloud trajectories instead of structural representations, we maintain spatial and temporal awareness of the states, enabling more precise garment folding. Additionally, employing generative model to create point cloud trajectories facilitates multi-category learning and generalization within a unified framework, yielding robust performance across diverse garment manipulation tasks.
@@
  • We introduce a novel framework for garment manipulation that disentangles state planning from action prediction. This hierarchical separation enables the model to focus on components, reducing training complexity.
@@
  • We developed the point cloud trajectory dataset for folding garments across multiple categories, accompanied by corresponding language descriptions.
@@ II-A Garment Folding
Garment folding, a challenging task owing to the deformable and high-dimensional nature of fabrics, has been extensively studied in the literature [12, 13, 14, 15].
@@
Cloth Funnels [8] trained robots to align garments into predefined deformable configurations, simplifying the manipulation process for downstream folding. UniGarmentManip [5] explored learning dense visual correspondences for garment manipulation with category-level generalization. FabricFlowNet [19] developed a cloth manipulation policy that leverages flow as both input and action representation to enhance performance. Zhou et al. [20] and Lin et al. [21] have proposed learning cloth dynamics to improve manipulation capabilities.
@@ II-B Language-Conditioned Object Manipulation
Language-conditioned manipulation has garnered significant attention due to its user-friendly interaction and strong generalization capabilities [23, 24, 25, 26, 27]. Stengel et al. [28] introduced a Transformer-based model for mapping natural language commands to robot actions in 3D tasks. CLIPort [29] combines CLIP’s semantic understanding with Transporter’s spatial precision for language-guided manipulation.
With recent advances in Large Language Models (LLMs), researchers have explored leveraging LLMs to integrate language understanding into cloth folding. GPT-Fabric [10] employs GPT to directly output actions, informing a robot where to grasp and pull a simple square-shaped fabric. Deng et al. [30] introduced a method for language-conditioned manipulation that utilizes graph dynamics to improve the understanding and prediction of deformable object behaviors.
@@ III Problem Formulation
The goal of language-guided garment folding is to generate a sequence of actions $\{\boldsymbol{a}_{i}\}_{i=1}^{n}$ to fold the garment into a target point cloud configuration $\mathcal{P}_{goal}$, given the 3D point cloud observation $\mathcal{P}\in\mathbb{R}^{N\times 3}$ with $N$ points and the language guidance $\mathcal{L}$.
@@
(1) Given the current $\mathcal{P}$ with $N$ points and $\mathcal{L}$, the goal of the point cloud trajectory generation model is to generate the trajectory $\mathcal{T}=\{\mathcal{P}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times N\times 3}$ that represents the evolution of the point cloud over time, where $M$ is the number of frames.
(2) Given two point clouds $(\mathcal{P},\mathcal{P^{\prime}})$, the aim of the ManiFoundation [11] model is to predict the action $\boldsymbol{a}=\{c_{i}\}_{i=1}^{n}$, which is defined as a set of contact syntheses. The contact synthesis for end effector $i$ is $c_{i}=(\boldsymbol{p},\boldsymbol{s})$, where $\boldsymbol{p}\in\mathbb{R}^{3}$ is the contact position and $\boldsymbol{s}\in\mathbb{R}^{3}$ is the corresponding motion direction based on the trajectory. We slice the generated garment point cloud trajectory $\mathcal{T}$ and input the segments into the ManiFoundation model to predict an action $\boldsymbol{a}=\mathcal{M}_{MF}(\mathcal{P},\mathcal{P^{\prime}})$, where $\{\mathcal{P},\mathcal{P^{\prime}}\}\subseteq\mathcal{T}$.
(3) Afterward, the robot executes the action $\boldsymbol{a}$, manipulating the garment to a new configuration $\mathcal{P}^{*}=\mathcal{M}_{Robot}(\mathcal{P},\boldsymbol{a})$.
@@ IV-A Overview
Our framework is illustrated in Figure 2. In this section, we provide a detailed explanation of language-guided point cloud trajectory generation (Sec. IV-B) and closed-loop manipulation (Sec. IV-C). Below, we first introduce the training and inference pipelines.
@@ IV-B2 Conditional Point Cloud Trajectory Generation
As discussed earlier, point cloud trajectories can model state transitions, allowing us to control intermediate states during garment folding. To achieve this, we design a point cloud trajectory generation model as a world model. This approach enables a comprehensive generation and control of garment movement over time, providing a complete understanding of its dynamics. From a spatial perspective, this method tracks all points in the initial frame and maps the subsequent trajectory of each point. Temporally, each frame represents the garment’s state at a specific moment.
@@
In training, both the spatial features of the point cloud $\mathcal{F}_{\mathcal{P}}$ and the language features $\mathcal{F}_{\mathcal{L}}$ are combined with the ground truth trajectory and used as conditional inputs to the CVAE encoder. This optimizes the distribution of the latent variable $z$. In both training and inference, samples are drawn from the distribution of $z$ and combined with the spatial features of the point cloud $\mathcal{F}_{\mathcal{P}}$. The language features $\mathcal{F}_{\mathcal{L}}$ are used as conditional inputs to the CVAE decoder to generate trajectory features $\mathcal{F}_{\mathcal{T}}\in\mathbb{R}^{M\times N\times 256}$. An additional MLP is used to reduce the dimensionality of the output, providing the generated trajectory $\mathcal{T}\in\mathbb{R}^{M\times N\times 3}$ for the current frame. The visualization of generated trajectories is in Figure 4.
@@ IV-B3 Semantic Generalization
We use a pre-trained large language model (Meta-Llama-3.1-8B-Instruct) [43] to extract semantic features. Specifically, we predefine a set of garment folding descriptions to strengthen the encoding process. During inference, we also employ LLaMA as the foundational language model to process user-provided guidance. The model matches user input to our predefined descriptions, effectively reducing the complexity of training on diverse language instructions by constraining conditions within a defined range. This approach leverages the large language model to achieve open-vocabulary generalization.
@@ IV-C2 ManiFoundation Model
Additionally, the predictions from the ManiFoundation model are influenced by random seeds. To mitigate this variability, we employ a model ensemble approach using 160 different random seeds to generate multiple predictions. When the distance between two prediction results is less than $\varepsilon$, they are considered similar and grouped. If a prediction result does not match any existing group, a new group is created. We compute the arithmetic mean of the point positions within the modal group and select the point closest to this average, along with its corresponding force, as the output contact point and force direction.
@@ V-C Baselines
For the garment folding task (Q1), we use UniGarmentManip [5] as the baseline. This method, previously demonstrated state-of-the-art (SOTA) performance in individual garment folding tasks [5], supports folding various garment types, thereby enabling a comprehensive comparison of folding performance across different categories. We use GPT-Fabric [10] as a baseline, which is a recent method that leverages large language models to select key points.
@@ V-D Results and Analysis
Although our approach is zero-shot on the CLOTH3D dataset, we achieve performance that matches or even surpasses the baseline, demonstrating the strong generalization of our method to a wide range of new garments. In garment folding, our model tends to fold clothing into a small and compact form, which aligns well with human expectations for neatness and compactness. In general, our approach outperforms the baselines in most tasks and metrics. Therefore, we can address Q1, demonstrating our superior performance in garment folding.
@@
Furthermore, our model supports folding based on user-provided language inputs, even if the folding sequence (Bottom $\rightarrow$ Left $\rightarrow$ Right) has not appeared in the training data. Figure 5 illustrates the folding results of our model under different language instructions.
@@ V-F Real-World Experiments
To obtain the garment point cloud, we first segmented the RGB image captured by an RGB-D camera using SAM2 [49] to generate a garment mask. This mask was then used to filter the corresponding depth data, extracting the actual point cloud of the clothing. The resulting points are then fed directly into our framework to forecast future folding states and the directions of contact-point movements. By executing these actions in sequence, the garment-folding process can be carried out in the real world. Because the sim-to-real gap in the point clouds is small, our approach can be transferred directly to real-world settings. The qualitative results are shown in Fig. 6, and the quantitative outcomes are listed in Table IV, with each garment tested over 10 trials.
TABLE IV: Real world results for our framework
No-sleeve Short-sleeve Long-sleeve Pants Rectangularity $\uparrow$ 0.94 0.91 0.87 0.85 Area Ratio $\downarrow$ 0.45 0.33 0.29 0.24 Success Rate $\uparrow$ 10/10 8/10 9/10 9/10
@@ VI Conclusion
In this work, we propose a comprehensive framework, MetaFold, for garment folding that supports multi-category folding tasks guided by user language instructions. This framework adopts a disentangled structure consisting of a point cloud trajectory generation model and a low-level action prediction model, utilizing closed-loop control for effective garment manipulation. We also constructed a point cloud trajectory dataset for garment folding, encompassing various folding methods across different garment types. Experimental results demonstrate that our approach achieves state-of-the-art performance in multi-category and language-guided garment folding tasks. We believe that MetaFold represents a significant step forward in applying trajectory generation to deformable object manipulation, marking an important milestone toward advancing spatial intelligence.
Acknowledgements
This research / project is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2024).